OpenStudy (anonymous):

If cos(20) - sin(20) = p, find the expression for cos(40) in terms of p.

OpenStudy (anonymous):

cos(40) = cos^2 (20) - sin^2 (20) = (cos20 - sin20)(cos20 + sin20) was my first thought.

OpenStudy (solomonzelman):

yes, but (cos20 + sin20) is going to unknown.

OpenStudy (solomonzelman):

You are thinking good though :)

OpenStudy (anonymous):

That's what the question is all about!

OpenStudy (solomonzelman):

Well, I guess you don't need more help with this one.

OpenStudy (anonymous):

I do :(

OpenStudy (anonymous):

@ganeshie8 @Kainui

OpenStudy (anonymous):

Oh nvm... sin(A/2) + cos(A/2) = sqrt(1 + sinA)

OpenStudy (lochana):

cos40 = (cos20 - sin20)(cos20 + sin20) what i'm suggesting is cos20 + sin20 = cos20 + cos70 = 2cos45cos(-25) = (cos25)sqrt2 --- 1 cos20 - sin20 =p (given) sin70 - sin20 = p 2sin45cos25=p (cos25)sqrt2 =p --- 2 using 1 and 2 ; cos20 + sin20 = p cos40 = p^2

OpenStudy (aum):

^^^^ If cos(20) - sin(20) = p can cos(20) + sin(20) also be p?

OpenStudy (aum):

\[ \cos(20) - \sin(20) = p \\ \text{Square both sides:} \\ \cos^2(20) + \sin^2(20) - 2\cos(20)\sin(20) = p^2 \\ 1 - 2\cos(20)\sin(20) = p^2 \\ 2\cos(20)\sin(20) = 1 - p^2\\ (\cos(20)+\sin(20))^2 = \cos^2(20) + \sin^2(20) + 2\cos(20)\sin(20) = 1 + 1 - p^2 \\ (\cos(20)+\sin(20)) = \sqrt{2-p^2} \\ \text{ } \\ \cos(40) = \cos^2 (20) - \sin^2 (20) = (\cos(20) - \sin(20))(\cos(20) + \sin(20)) = p\sqrt{2-p^2} \\ \]